# Adapting In-Domain Few-Shot Segmentation to New Domains without Retraining

### v2

@@ 3.4 Informative Structure Adaptation
We incorporate the proposed ISI and PSA modules into the Informative Structure Adaptation (ISA) method, as shown in Algorithm 1. Specifically, we first use ISI to select the trainable parameters $\theta_{\texttt{tr}}$ for test-time training. Then, we use PSA with hierarchically constructed training pairs to train the selected model parameters by gradually increasing the number of support shots $n$ from $1$ to $K-1$:
@@
Our ISA method can also be applied to transformer-based methods [67, 66] by adaptively selecting and tuning the fully-connected layers of the backbone during inference.
@@ 4 Experiments
We adopt the popular few-shot semantic segmentation model SSP [11] as our baseline, trained on Pascal VOC [9] source domain dataset. We directly apply our method to the public released, well-trained SSP model with a ResNet-50 [15] backbone, without any re-training on the source domain dataset. For test-time training, we use the SGD optimizer with a learning rate of 1e-3 and one training iteration to update the trainable model parameters. Following previous works [25, 45], we evaluate all methods on four datasets with distinct domain shifts: Deepglobe [7] for satellite images with seven categories, ISIC2018 [6, 51] for medical images with three types of skin lesions, Chest X-Ray [2, 21] for medical screening images, and FSS-1000 [27] for 1000-class daily objects. The input images are resized to $400\times 400$ pixels. In the 1-shot setting, we apply data augmentation to support images to generate two additional support images for test-time training. We use the mean Intersection-over-Union (mIoU) for evaluation. All experiments are conducted on a Tesla V100 GPU.
@@
Table 3: Results of applying our method to FSS/CD-FSS models with ResNet and Transformer backbones. “In Domain?” indicates whether the model is designed for in-domain FSS. “w/ Ours” denotes that the model is equipped with our ISA method. “$\Delta$” represents the performance improvement. All models are trained on Pascal VOC dataset (except for PerSAM), and evaluated on four CD-FSS datasets. The average 5-shot mIoU performance is reported.
@@
SSP [11] ResNet-50 63.0 70.3 +7.3
PANet [54] ResNet-50 55.1 61.8 +6.7
FPTrans [66] Transformer 60.3 67.4 +7.1
PerSAM [67] Transformer 61.5 67.9 +6.4
DR-Adapter [45] ResNet-50 65.4 70.9 +5.5
Figure 3: Selected trainable layer distribution of the informative structure identification (ISI) module.
@@ 4.1 Comparison with State-of-the-Arts
In Table 1, we compare our method with existing cross-domain few-shot semantic segmentation methods. Our method substantially outperforms the baseline method SSP [11], with a 4.1/7.3 mIoU average improvement in the 1-shot/5-shot settings across all datasets. Additionally, our method surpasses previous CD-FSS SOTA methods, PMNet [3], ABCDFSS [18], APSeg [17] and DR-Adapter [45], by a large margin in the 5-shot setting. APSeg achieves slightly better performance than our method (61.3 $v.s.$ 60.9), primarily because their ViT backbone is more powerful than our ResNet-50 backbone. Note that all other CD-FSS methods require extensive retraining on the source domain dataset to learn transferable, domain-agnostic features for domain generalization. In contrast, our method can effectively and efficiently adapt existing well-trained FSS models for segmenting objects of novel classes under domain shifts without any retraining.
To further validate the effectiveness of our method, we follow RD [57] to evaluate our method on the SUIM dataset. All models are trained on Pascal VOC dataset and evaluated on SUIM [20] dataset. Table 2 shows that our method improves the SOTA performance from 40.3 to 44.1 mIoU, beating other popular methods, including ASGNet [26], HSNet [36], SCL [62], RD [57], DAM [4], MMT [58] and DR-Adapter [45].
Recently, IFA [38] sets a new SOTA on CD-FSS benchmarks, but they adopt a distinct evaluation protocol. For a fair comparison, we adopt their code and evaluation protocol to implement and evaluate our method. When combined with their baseline model, our method achieves 72.8 mIoU, surpassing their reported 71.4 mIoU.
@@
Generalized to Other Methods. Our ISA method is general and can be applied to other FSS methods to address cross-domain few-shot semantic segmentation. As shown in Table 3, when equipped with our ISA method, two ResNet-based in-domain FSS methods, SSP [11] and PANet [54] achieve substantially better performance on CD-FSS. Our method can improve transformer-based in-domain FSS method FPTrans [66], boosting the performance to 67.4 mIoU for CD-FSS. Our method can also adapt the powerful SAM-based method PerSAM [67] for CD-FSS, achieving remarkable 67.9 mIoU. Our method can further improve existing CD-FSS methods, as evidenced by the 5.5 mIoU improvement on DR-Adapter [45] when combined with our ISA method.
@@
Mixstyle [69] DSU [28] NP [12] TTT [49] Tent [52]
@@
Adapter [5] BitFit [61] PerSAM [67] Matcher [34] Ours
@@
Comparison with More Related Methods. Table 4 compares our method with domain generalization (Mixstyle [69], DSU [28], and NP [12]), test-time training (TTT [49] and Tent [52]), PEFT-based methods (Adapter [5] and BitFit [61], PEFT for Parameter-Efficient Fine-Tuning), and SAM-based methods (PerSAM [67] and Matcher [34]) to further demonstrate the superiority of our approach. Our method significantly outperforms other related methods.
@@ 4.2 Informative Structure Adaptation Analysis
Table 5: Results of ablation studies for the self-guiding test-time training method. “MSA” denotes the model structure adaptation baseline, “ISI” denotes the informative structure identification module, and “PSA” denotes the progressive structure adaptation module.
@@
Table 7: Results of using various progressive training strategies in the progressive structure adaptation (PSA) module. The “1/2/3/4” denotes training models using HTC-constructed samples with 1/2/3/4 support shots. The $\rightarrow$ denotes the sequential training procedure.
1 2 3 4 1$\rightarrow$4 2$\rightarrow$4 3$\rightarrow$4
@@
1$\rightarrow$2$\rightarrow$4 1$\rightarrow$3$\rightarrow$4 2$\rightarrow$3$\rightarrow$4 1$\rightarrow$2$\rightarrow$3$\rightarrow$4
@@
Discussions on Foundation Model-based Methods. Foundation models, such as SAM [23] and CLIP [40], are typically trained on large-scale web-collected data, resulting in excellent generalization on natural images. However, due to data domain limitations, they often underperform in unseen domains like medical images, remote sensing images, or industrial images. Additionally, foundation models are typically built on large backbone models, leading to computation, deployment, and storage challenges. In contrast, our method is specifically designed to address generalization to novel classes and unseen domains, featuring lightweight computation, a simple network architecture, few model parameters, and easy deployment. Our method is general and can flexibly equip foundation models to address domain shifts, evidenced in Table 3.
@@ 5 Conclusion
We propose, for the first time, a novel framework Informative Structure Adaptation (ISA) that generalizes well-trained in-domain few-shot segmentation models to new domains without retraining. Our Informative Structure Identification (ISI) adaptively identifies domain-specific model structures by measuring parameter importance with a novel structure Fisher score. Furthermore, we propose the Progressive Structure Adaptation (PSA) to progressively adapt the selected informative model structures during inference, utilizing hierarchically constructed training samples with an increasing number of support shots. The ISA method combines these strategies to effectively address domain shifts in CD-FSS, and equips existing few-shot segmentation models with flexible adaptation capabilities for new domains.

### v3

@@ 3.4 Informative Structure Adaptation
We incorporate the proposed ISI and PSA modules into the Informative Structure Adaptation (ISA) method, as shown in Algorithm 1. Specifically, we first use ISI to select the trainable parameters $\theta_{\texttt{tr}}$ for test-time training. Then, we use PSA with hierarchically constructed training pairs to train the selected model parameters by gradually increasing the number of support shots $n$ from $1$ to K1K-1:
@@
Our ISA method can also be applied to transformer-based methods [zhang2023personalize, zhang2022feature] by adaptively selecting and tuning the fully-connected layers of the backbone during inference.
@@ 4 Experiments
We adopt the popular few-shot semantic segmentation model SSP [fan2022self] as our baseline, trained on Pascal VOC [everingham2010pascal] source domain dataset. We directly apply our method to the public released, well-trained SSP model with a ResNet-50 [he2016deep] backbone, without any re-training on the source domain dataset. For test-time training, we use the SGD optimizer with a learning rate of 1e-3 and one training iteration to update the trainable model parameters. Following previous works [lei2022cross, su2024domain], we evaluate all methods on four datasets with distinct domain shifts: Deepglobe [demir2018deepglobe] for satellite images with seven categories, ISIC2018 [codella2019skin, tschandl2018ham10000] for medical images with three types of skin lesions, Chest X-Ray [candemir2013lung, jaeger2013automatic] for medical screening images, and FSS-1000 [li2020fss] for 1000-class daily objects. The input images are resized to $400\times 400$ pixels. In the 1-shot setting, we apply data augmentation to support images to generate two additional support images for test-time training. We use the mean Intersection-over-Union (mIoU) for evaluation. All experiments are conducted on a Tesla V100 GPU.
@@
Table 3: Results of applying our method to FSS/CD-FSS models with ResNet and Transformer backbones. “In Domain?” indicates whether the model is designed for in-domain FSS. “w/ Ours” denotes that the model is equipped with our ISA method. $\Delta$” represents the performance improvement. All models are trained on Pascal VOC (except for PerSAM), and evaluated on four CD-FSS datasets. The average 5-shot mIoU performance is reported.
@@
SSP [fan2022self] ResNet-50 63.0 70.3 +7.3
PANet [wang2019panet] ResNet-50 55.1 61.8 +6.7
FPTrans [zhang2022feature] Transformer 60.3 67.4 +7.1
PerSAM [zhang2023personalize] Transformer 61.5 67.9 +6.4
DR-Adapter [su2024domain] ResNet-50 65.4 70.9 +5.5
Figure 3: Selected trainable layer distribution of the informative structure identification (ISI) module. The “selected ratio” denotes the frequency with which each layer is selected across the entire dataset.
@@ 4.1 Comparison with State-of-the-Arts
In Table 1, we compare our method with existing cross-domain few-shot semantic segmentation methods. Our method substantially outperforms the baseline method SSP [fan2022self], with a 4.1/7.3 mIoU average improvement in the 1-shot/5-shot settings across all datasets. Additionally, our method surpasses previous CD-FSS SOTA methods, PMNet [chen2024pixel], ABCDFSS [herzog2024adapt], APSeg [he2024apseg] DR-Adapter [su2024domain] and DMTNet [chen2024cross], by a large margin in the 5-shot setting. APSeg achieves slightly better performance than our method (61.3 $v.s.$ 60.9), primarily because their ViT backbone is more powerful than our ResNet-50 backbone. Note that most other CD-FSS methods (except for ABCDFSS) require extensive source domain retraining, typically dozens of hours, on the source domain dataset to learn transferable, domain-agnostic features for domain generalization. In contrast, our method can effectively and efficiently adapt existing well-trained FSS models for segmenting objects of novel classes under domain shifts without any source domain retraining.
To further validate the effectiveness of our method, we follow RD [wang2022remember] to evaluate our method on the SUIM dataset. All models are trained on Pascal VOC dataset and evaluated on SUIM [islam2020semantic] dataset. Table 2 shows that our method improves the SOTA performance from 40.3 to 44.1 mIoU, beating other popular methods, including ASGNet [li2021adaptive], HSNet [min2021hypercorrelation], SCL [zhang2021self], RD [wang2022remember], DAM [chen2024dense], MMT [wang2023mmt] and DR-Adapter [su2024domain].
Recently, IFA [nie2024cross] sets a new SOTA on CD-FSS benchmarks, but they adopt a distinct evaluation protocol. For a fair comparison, we adopt their code and evaluation protocol to implement and evaluate our method. When combined with their baseline model, our method achieves 72.8 mIoU, surpassing their reported 71.4 mIoU.
@@
Generalized to Other Methods. Our ISA method is general and can be applied to other FSS methods to address cross-domain few-shot semantic segmentation. As shown in Table 3, when equipped with our ISA method, two ResNet-based in-domain FSS methods, SSP [fan2022self] and PANet [wang2019panet] achieve substantially better performance on CD-FSS. Our method can improve transformer-based in-domain FSS method FPTrans [zhang2022feature], boosting the performance to 67.4 mIoU for CD-FSS. Our method can also adapt the powerful SAM-based method PerSAM [zhang2023personalize] for CD-FSS, achieving remarkable 67.9 mIoU. Our method can further improve existing CD-FSS methods, as evidenced by the 5.5 mIoU improvement on DR-Adapter [su2024domain] when combined with our ISA method.
@@
Mixstyle [zhou2021domain] DSU [li2022uncertainty] NP [fan2022towards] TTT [sun2020test] Tent [wang2020tent]
@@
Adapter [chen2022adaptformer] BitFit [zaken2022bitfit] PerSAM [zhang2023personalize] Matcher [liu2023matcher] Ours
@@
Comparison with More Related Methods. Table 4 compares our method with domain generalization (Mixstyle [zhou2021domain], DSU [li2022uncertainty], and NP [fan2022towards]), test-time training (TTT [sun2020test] and Tent [wang2020tent]), PEFT-based methods (Adapter [chen2022adaptformer] and BitFit [zaken2022bitfit], PEFT for Parameter-Efficient Fine-Tuning), and SAM-based methods (PerSAM [zhang2023personalize] and Matcher [liu2023matcher]) to further demonstrate the superiority of our approach. Our method significantly outperforms other related methods.
@@ 4.2 Informative Structure Adaptation Analysis
Table 5: Results of ablation studies for our method. “MSA” denotes the model structure adaptation baseline, “ISI” denotes the informative structure identification module, and “PSA” denotes the progressive structure adaptation module.
@@
Table 7: Results of using various training strategies in the progressive structure adaptation (PSA) module. The “1/2/3/4” denotes training models using HTC-constructed samples with 1/2/3/4 support shots. The $\rightarrow$ denotes the sequential training procedure.
[…]
@@
[…]
@@
Discussions on Foundation Model-based Methods. Foundation models, such as SAM [kirillov2023segment] and CLIP [radford2021learning], are typically trained on large-scale web-collected data, resulting in excellent generalization on natural images. However, due to data domain limitations, they often underperform in unseen domains like medical images, remote sensing images, or industrial images. Additionally, foundation models are typically built on large backbone models, leading to computation, deployment, and storage challenges. In contrast, our method is specifically designed to address generalization to novel classes and unseen domains, featuring lightweight computation, a simple network architecture, few model parameters, and easy deployment. Our method is general and can flexibly equip foundation models to address domain shifts, evidenced in Table 3.
@@ 5 Conclusion
We propose, for the first time, a novel framework Informative Structure Adaptation (ISA) that generalizes well-trained in-domain few-shot segmentation models to new domains without source domain retraining. Our Informative Structure Identification (ISI) adaptively identifies domain-specific model structures by measuring parameter importance with a novel structure Fisher score. Furthermore, we propose the Progressive Structure Adaptation (PSA) to progressively adapt the selected informative model structures during inference, utilizing hierarchically constructed training samples with an increasing number of support shots. The ISA method combines these strategies to effectively address domain shifts in CD-FSS, and equips existing few-shot segmentation models with flexible adaptation capabilities for new domains.
Acknowledgements
This work is supported in part by the National Natural Science Foundation of China (62192783, 62276128, 62406140), Young Elite Scientists Sponsorship Program by China Association for Science and Technology (2023QNRC001), the Key Research and Development Program of Jiangsu Province under Grant (BE2023019) and Jiangsu Natural Science Foundation under Grant (BK20221441, BK20241200)
6 Discussion on Fisher Information for ISA
We provide a detailed explanation of the rationale for using Fisher information (FI) as a criterion for selecting and tuning layers and parameters in our ISA method. We hope this discussion helps explain why Fisher information is particularly suitable for CD-FSS tasks.
6.1 Theoretical Justification
When considering the importance of parameters, we generally aim to understand how much changing a parameter will impact the model’s output. Let us denote $p_{\theta}(y|x)$ as the output distribution over $y$ produced by a model with parameter vector $\theta\in\mathbb{R}^{|\theta|}$, given input $x$.
One way to measure the impact of a change in parameters on the model’s prediction is by computing the Kullback-Leibler divergence (KL divergence) between the output distributions:
$\text{D}_{\text{KL}}(p_{\theta}(y|x)||p_{\theta+\delta}(y|x)),$ (8)
where $\delta\in\mathbb{R}^{|\theta|}$ represents a small perturbation. It can be shown that as $\delta\to 0$, we can perform a second-order Taylor expansion at $\theta$, yielding:
$\mathbb{E}_{x}\left[\text{D}_{\text{KL}}(p_{\theta}(y|x)||p_{\theta+\delta}(y|x))\right]=\delta^{T}F_{\theta}\delta+O(\delta^{3}),$ (9)
where $F_{\theta}\in\mathbb{R}^{|\theta|\times|\theta|}$ is the Fisher information matrix, which is defined as:
$F_{\theta}=\mathbb{E}_{x\sim p(x)}\mathbb{E}_{y\sim p_{\theta}(y|x)}\left[\nabla_{\theta}\log p_{\theta}(y|x)\nabla_{\theta}\log p_{\theta}(y|x)^{T}\right].$ (10)
Due to the large size of $F_{\theta}$, it is computationally intractable to compute it directly. As a result, we typically ignore the interactions between different parameters, and focus on the importance of each individual parameter in isolation. Therefore, following prior work, we approximate $F_{\theta}$ by considering only its diagonal, which can be represented as a vector in $\mathbb{R}^{|\theta|}$.
Thus, we obtain the following approximation for the Fisher information:
$\hat{F}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{y\sim p_{\theta}(y|x_{i})}\left[\left(\nabla_{\theta}\log p_{\theta}(y|x_{i})\right)^{2}\right],$ (11)
where $\hat{F}_{\theta}\in\mathbb{R}^{|\theta|}$.
Although the Fisher information requires taking an expectation over the output distribution, in a supervised learning setting, we can leverage the availability of the ground-truth label $y_{i}$ for each sample $x_{i}$. This allows us to replace the expectation $\mathbb{E}_{y\sim p_{\theta}(y|x_{i})}\left[\left(\nabla_{\theta}\log p_{\theta}(y|x_{i})\right)^{2}\right]$ in Eq. (1.4) with the squared gradient corresponding to the true label, $\left(\nabla_{\theta}\log p_{\theta}(y_{i}|x_{i})\right)^{2}$.
This approach, known as ”Empirical Fisher,” provides a straightforward method for estimating the parameter importance heuristically. It can be computed as follows:
$\hat{F}_{\theta}\approx\frac{1}{N}\sum_{i=1}^{N}\left(\nabla_{\theta}\log p_{\theta}(y_{i}|x_{i})\right)^{2},$ (12)
where $p_{\theta}(y_{i}|x_{i})$ denotes the output probability of $y_{i}$ given input $x_{i}$ and parameters $\theta$, $N$ is the number of samples, and $\hat{F}_{\theta}\in\mathbb{R}^{|\theta|}$.
6.2 Intuitive Interpretation
Each element of the approximated Fisher Information vector $\hat{F}_{\theta}$ represents the expected squared gradient of the log-likelihood function with respect to a given parameter. If a particular parameter strongly influences the model’s output, its corresponding entry in $\hat{F}_{\theta}$ will be large. Therefore, we can reasonably interpret $\hat{F}_{\theta}$ as an approximation of the importance of each parameter in determining the model’s output.
We hope this discussion provides a better understanding of the theoretical foundation and the intuitive interpretation of using Fisher information for selecting and tuning parameters.
6.3 More Discussions
The primary advantage of Fisher Information over other methods [chen2018coupled, soen2021variance] lies in its principled, information-theoretic approach to quantifying the importance of model parameters. Unlike heuristic methods, which rely on rules of thumb or assumptions about parameter relevance, Fisher Information provides a mathematically rigorous measure of how sensitive the model’s output is to parameter changes. This makes FI particularly effective in settings with limited data, such as CD-FSS, where traditional methods like heuristic or gradient-based approaches may fail due to instability or high noise levels. For instance, heuristic methods tend to be inconsistent in few-shot learning settings, and gradient-based methods, while effective in large-data settings, can be noisy and less reliable in few-shot regimes due to data scarcity. On the other hand, Fisher Information provides a more stable and consistent criterion for identifying parameter importance, which helps prioritize the parameters most critical to the model’s performance.
In our work, we focus on model structure adaptation, where we leverage Fisher Information to identify and fine-tune only the most critical layers of the model, instead of fine-tuning all parameters. This approach is based on our observation that Fisher Information is sparsely distributed across the model, with many convolutional layers exhibiting low-value entries. To address this, we introduce structure Fisher scores, which compute the maximum Fisher Information across all parameters within each layer. Layers with higher structure Fisher scores are typically more influential in the optimization process and, thus, more important for model adaptation. By focusing on the most informative layers, we ensure that fine-tuning efforts are more efficient, preserving the model’s ability to generalize while reducing the risk of overfitting in few-shot learning scenarios.
Why is Fisher Information particularly suitable for solving the problem of CD-FSS? The key challenge in cross-domain few-shot segmentation is the risk of overfitting, due to the limited number of training samples from the target domain. Fisher Information is particularly suited for this problem because it helps identify the model layers that are most crucial for generalization across domains. By targeting only the layers with the highest Fisher scores for adaptation, we can reduce the likelihood of overfitting, ensuring that the model focuses on the most important features for domain transfer. This targeted adaptation prevents unnecessary fine-tuning of less important parameters, which could introduce noise or instability, especially with limited data. Consequently, the model can better generalize to unseen domains without compromising performance.
In summary, Fisher Information offers a more stable, principled, and efficient approach to parameter selection and model adaptation compared to traditional heuristic or gradient-based methods. It is particularly well-suited for cross-domain few-shot segmentation, as it enables selective and targeted adaptation, helping to mitigate overfitting while enhancing generalization across domains.
7 Limitations
Although the proposed method demonstrates substantial performance improvements, it has inherent limitations that should be considered. The additional modules introduced in the approach add complexity, potentially reducing processing speed due to extra model forwarding steps and increased computational overhead. As a result, while the method excels in performance, it may not be suitable for applications that require high-speed processing, such as real-time systems or tasks with strict time constraints.
Despite these challenges, the proposed method is well-suited for performance-demanding applications where accuracy takes precedence over speed. Future work could explore optimizations to balance the trade-off between performance and computational efficiency, possibly through more efficient model design or advanced hardware acceleration techniques. These improvements could make the method more versatile for a broader range of practical applications.
We further propose a fast ISA method based on our experimental analysis. Specifically, we replace ISI by directly selecting the “layer3.5.conv1” and the last convolutional layers as the trainable layers. Additionally, we replace PSA with a ($2\rightarrow 3\rightarrow 4$)-based progressive training strategy, and randomly select only one training pair for each pseudo query data. The proposed fast ISA method achieves 70.0 mIoU, with a considerable improvement on the running speed and a marginal performance drop compared with the original ISA method.